OpenStudy (anonymous):

FOR REDUCTION OF ORDER (D^2+4)y=sinx @TuringTest :)) i have the answer but i dont know how it came with it :)

OpenStudy (wasiqss):

well i can tell you how to do it

OpenStudy (anonymous):

that's great! can you help me? please? :)

OpenStudy (wasiqss):

do you want complimentary or particular solution

OpenStudy (anonymous):

both? and the general solution? :)))) please?

OpenStudy (wasiqss):

y complimentary =d^2=-4 d=+2i, -2i hence Y complimentary =C1 cos2x+C2 Sin2x

OpenStudy (wasiqss):

Y particular = (Sinx)/(D^2 +4) now the cofficient of x (that is with sin =1) hence we put the square of the negative value of coffiecient of x in the variable D hence it becomes Sinx/(- (1^2)+4) hence Y particular=Sinx/3

OpenStudy (anonymous):

can you give me a detailed solution? like in integrating it? or what? does it need wronkian in reduction?

OpenStudy (wasiqss):

Lol its the solution and detailed one and the general solution is Yparticular + Y general

OpenStudy (anonymous):

so you mean the reduction of order is a simplier method compared to variation? or it depends?

OpenStudy (wasiqss):

well i did the differential equations this way and i find it simple

OpenStudy (wasiqss):

u in which grade

OpenStudy (anonymous):

2nd year college

OpenStudy (wasiqss):

im a first year college student though :D

OpenStudy (anonymous):

hmm.. i thinkn you have k12 :)

OpenStudy (anonymous):

so what's the steps that im going to put on my paper? :) first the yc then the yp immediately? and lastly the y?

OpenStudy (wasiqss):

yehh exactly :)

OpenStudy (wasiqss):

what is K12

OpenStudy (anonymous):

woah. what a short process?

OpenStudy (wasiqss):

yehh i will be glad to help u in these type of questions :)_

OpenStudy (anonymous):

hence we put the square of the negative value of coffiecient of x in the variable D hence it becomes ??? wat do u mean of that? can u clarify it with me?

OpenStudy (wasiqss):

like if its Sin 2x then the cofficient of x=2 and hence we would put the square that is 4 in this case and negative means i will add - to the square of the coffiecient .

OpenStudy (wasiqss):

but remember we can only do that when it is D^2

OpenStudy (anonymous):

so i shall only follow the process that you gave to me?

OpenStudy (wasiqss):

yehh u should

OpenStudy (anonymous):

Y particular = (Sinx)/(D^2 +4) now the cofficient of x (that is with sin =1) hence we put the square of the negative value of coffiecient of x in the variable D hence it becomes Sinx/(- (1^2)+4) hence Y particular=Sinx/3 so i will put this on my paper? lol

OpenStudy (anonymous):

yp= sinx / (d^2+4)

OpenStudy (wasiqss):

Yparticular = (Sinx)/(D^2 +4) now the cofficient of x (that is with sin =1) hence we put the square of the negative value of coffiecient of x in the variable D^2 hence it becomes Sinx/(- (1^2)+4) hence Y particular=Sinx/3

OpenStudy (wasiqss):

now thats perfect

OpenStudy (anonymous):

now the cofficient of x (that is with sin =1)???? this?

OpenStudy (wasiqss):

yehh

OpenStudy (anonymous):

wat do u mean? is it constant? sin =1?

OpenStudy (wasiqss):

i mean if its sin3x then constant cofficient=3

OpenStudy (anonymous):

ahh.. now i know! i get it. lol HAHA so if sin3x the it could sin3x/-3^2+4? i jst change 1 to 3

OpenStudy (wasiqss):

yayyy finally u got it :)

OpenStudy (anonymous):

hmm.. so the short process of yours will give me high score? :))

OpenStudy (wasiqss):

yehhh :)

OpenStudy (anonymous):

HAHA. thanks friend. i hope i can count on you next time :)

OpenStudy (wasiqss):

Lol probability is low though :D

OpenStudy (anonymous):

HAHA.. lol :)

OpenStudy (turingtest):

sorry, I need to review reduction of order...

OpenStudy (anonymous):

hey turing test. i have a question. ahm. can wolfram able to answer laplace problems?

OpenStudy (amistre64):

yes it can

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=laplace+inverse